OpenStudy (anonymous):

find the indicated outputs for f(x)=3x^2-3x answer f(0)=? f(-1)=? f(2)=?

OpenStudy (anonymous):

uhm... replace Plop

OpenStudy (anonymous):

Yeah. Replace x with the numbers indicated in the functional brackets. Do basic calculator math. Get answers.

OpenStudy (anonymous):

f(0)=0 f(-1)=0 f(2)= 6

OpenStudy (anonymous):

Yes!

OpenStudy (phi):

redo f(-1): f(x)=3x^2-3x = 3* (-1*-1) - 3* -1 = 3*1 + 3= 6

OpenStudy (anonymous):

Woops. You're right, teh first term is x^2, not x^3

OpenStudy (anonymous):

I can't help people making mistakes like this. :c

OpenStudy (phi):

they can always ask if it doesn't make sense. It might even help them.